# Federated Variational Inference for Bayesian Mixture Models

### v1

@@ 4 Federated Variational Inference
In this section, we propose a scalable clustering algorithm for large categorical datasets, FedMerDel (Federated MerDel), which addresses the challenge of large $N$ and enables federated learning across multiple nodes.
@@
The second ‘global’ step then combines the local clustering structures. Local clusters are frozen, and we merge similar local clusters into ‘global’ clusters. This is conceptually similar to approaches used in SNOB (Zuanetti et al., 2018) and SIGN (Ni et al., 2019), but both use MCMC methods which are computationally costly; local clustering in SNOB takes over 4 hours for a batch of size $N=1000$ in the authors’ simulations. We utilise a version of the variational merge move seen in MerDel to implement a ‘global merge’. By freezing local clusters in the global merge, the proposed clustering structure and ‘global ELBO’ across all observations can be efficiently calculated by utilising stored parameter values. The variational framework allows us to accept or reject global merges in a principled manner. Importantly, the full dataset is not required in the global merge.
In contrast to the approach considered in Section 3.2, instead of optimising individual components of the existing classical variational algorithm, our approach is designed to present a scalable clustering model which avoids these bottlenecks entirely by adopting a distributed computing approach in which each data batch is analysed independently and in parallel before ‘global merge’ steps are used to obtain a (global) mixture model for the full dataset. This federated learning approach eliminates the need for frequent communication and synchronization during the main variational updates, while data remains decentralised.
@@ 4.1 Estimating Global Clusters
Assume our data has been split into $B$ batches and MerDel has been run on each batch, with $K_{b}$ local clusters in Batch $b$. Let $K=\sum_{b=1}^{B}K_{b}$ be the total number of local clusters. We assume we are in the setting with no variable selection. In the global step, we first ‘combine’ all our existing variational parameters from each batch. For $q(\pi)$, we concatenate all our $K_{b}$-length $\alpha$ vectors to one $K$-length $\alpha$ vector. Each $q(\phi_{kj})$ is already independent for each local cluster $k$ (and each covariate $j$). We create a new matrix of responsibilities $r_{nk}$ as an $N\times K$ block diagonal matrix, consisting of blocks of $N_{b}\times K_{b}$ responsibility matrices from each batch. $\sum_{k=1}^{K}r_{nk}=1$ still holds for all $n$.
Let clusters $k_{1}$ and $k_{2}$ be considered for a merge. As in the local merge, we reassign responsibilities to cluster $k_{1}$ via $r^{new}_{nk_{1}}=r_{nk_{1}}+r_{nk_{2}}$ for all $n=1,...,N$. This can be viewed as summing columns $k_{1}$ and $k_{2}$ in our responsibility matrix. We then use the definitions of $\alpha^{\ast}$, $\epsilon^{\ast}$ (Equations (10), (11)) as follows to update cluster-specific parameters:
$\alpha^{new}_{k_{1}}$ $=\alpha_{0}+\sum_{n=1}^{N}(r_{nk_{1}}+r_{nk_{2}})$ (12)
$=\alpha^{\ast}_{k_{1}}+\alpha^{\ast}_{k_{2}}-\alpha_{0}$ (13)
$\epsilon^{new}_{k_{1}jl}$ $=\epsilon_{jl}+\sum_{n=1}^{N}\mathbb{I}(x_{nj}=l)(r_{nk_{1}}+r_{nk_{2}})$ (14)
$=\epsilon^{\ast}_{k_{1}jl}+\epsilon^{\ast}_{k_{2}jl}-\epsilon_{j}$ (15)
This avoids calculating sums over $N$. To complete the merge, we remove all parameters associated with cluster $k_{2}$ $\alpha^{\ast}_{k_{2}}$, $\epsilon^{\ast}_{k_{2}j}$ and $r_{nk_{2}}$ for all $n=1,...,N$, $j=1,...,P$. Unlike the variational merge moves, we do not perform an extra variational E/M step to allow observations to move in and out of the new cluster. This would require a whole-dataset update of responsibilities $r_{nk}$. The efficiency of the global merge hinges on the fact that local clusters are frozen. We propose global merge candidates in two ways; a greedy search and a random search.
Greedy Search
We prevent merges between two clusters from the same batch with a greedy search, on the hypothesis that such merges would already have taken place at the local stage. This can be carried out in a sequential manner by examining the correlation/divergences between Cluster 1 of Batch 1 and all clusters in Batch 2. Once Cluster 1 has merged with any cluster in Batch 2, we move to Batch 3 and consider merging with the clusters in this batch. We repeat for all batches, before moving to Cluster 2 of Batch 1 and so on until we have gone through all clusters in Batch 1. We then repeat the process with Batches 2, 3, …, B, looking only at Batches $b+1,...,B$ for Batch $b$.
Random Search
Similarly to local merge moves, we calculate correlations/divergences between all $K$ clusters, and pick the merge candidate pair randomly between the 3 pairs with the highest correlation (provided correlation $>$ 0.05) or lowest divergence. There is no restriction on proposing merges between two clusters of the same batch. This stops after a fixed number of global merge moves (e.g. 10) have been rejected consecutively.
@@ 4.2 Efficient Global ELBO Calculation
The ELBO variational objective function allows us to accept or reject the global merge under a rigorous probabilistic framework. For the global merge to be accepted, the ELBO must be higher in the new, merged model. As we now consider the full dataset, many terms in the overall ELBO function involve all observations and clusters. ‘Freezing’ local clusters allows us to efficiently calculate the majority of the ELBO terms while enabling the model to handle large-scale data. Two sums involving N in the full ELBO function can be written as a sum over smaller $K$, $P$, $L_{j}$ of sufficient statistics which can in turn be calculated as a linear function of previously calculated $\alpha^{\ast}$ and $\epsilon^{\ast}$: $T_{k}=\sum_{n=1}^{N}r_{nk}=\alpha^{\ast}_{k}-\alpha_{0}$ and $S_{kjl}=\sum_{n=1}^{N}r_{nk}\mathbb{I}(x_{nj}=l)=\epsilon^{\ast}_{kjl}-\epsilon_{jl}$. This vastly reduces the number of operations required.
The final term involving $N$ is an assignment entropy term, $\sum_{n=1}^{N}\sum_{k=1}^{K}r_{nk}\ln r_{nk}$, which is not linear in any previously calculated parameters. This is additive over $N$ and $K$, so we can calculate this value separately for each batch (in parallel if required) before adding all results together. However, in our greedy search where we never combine two clusters from the same batch, $r_{nk_{1}}$ and/or $r_{nk_{2}}$ will always be zero for any merge, for all $n=1,...,N$. Observation $n$ is only explained by clusters originating from certain batches of data and a merge is only proposed from a new batch where $n$ is not explained by any clusters associated with that batch. This allows greedy search to be faster, although merges could potentially be missed.
Full details of the global ELBO calculation can be found in Appendix F, where all terms in the ELBO can be written without $X$. Crucially, the global merge move allows us to seek a clustering structure for the full dataset without any scan over the full data matrix $X$, which is essential for federated learning. Data from separate entities remains decentralised, where only statistics from each batch are required to be shared, allowing for increased data security and privacy. We also note that parallelisation in local batches, as in Section 3.2, could still be applied to FedMerDel for additional gains, given sufficient computational resources.
4.3 Implementation
The model builds upon an existing R package, VICatMix (Rao & Kirk, 2024), which applies VI for binary and categorical Bayesian finite mixture models (with no merge/delete moves). Rcpp and RcppArmadillo are used to accelerate computation with C++. MerDel runs on a single-core processor, where, for example, inference for a mixture model fitted to a binary dataset of size $N=2000$, $P=100$ generally converges in less than a minute.
FedMerDel allows MerDel to be run independently on a different core per batch of data; a subsequent global merge takes only a few seconds at most in all simulations. Reported ‘wall-clock times’ take this parallelisation into account where clearly stated. Simulations were all run on a university multi-core HPC cluster.
5 Examples
@@ 5.1 Simulation Study
In this section, all simulated data is binary, with the data generating mechanism detailed in the Appendix. Additional simulation studies comparing merge criteria, looking at scenarios with simulated categorical data with 3+ categories, and looking at variable selection can also be found in the Appendix.
Frequency of Merge/Delete Moves
@@
In Section 3.1, we described the trade-off between more or fewer merge/delete moves in MerDel. We ran simulation studies comparing different values of the parameter ‘laps’ representing the frequency of merge/delete moves in a run of MerDel, with laps $\in\{1,2,5,10\}$. We also compared to a model with just merge/delete moves after performing one ‘E’ and ‘M’ step after initialisation (‘laps = 0’), and the variational model with no merge/delete moves (‘laps = 10000’). We used correlation for merge criteria.
We compared computational efficiency by measuring the wall-clock time taken for convergence. We additionally compared the number of non-empty clusters found, and the Adjusted Rand Index (ARI) with the simulated ‘true’ labels. Details of the simulated datasets are in Appendix I.3.
Global Merge Simulations
@@
We evaluated the clustering performance when splitting data into batches and performing a global merge as detailed in Section 4, and compared this with the performances of MerDel applied to the full dataset, and parallelised MerDel as described in Section 3.2. We looked at binary data with sizes varying between $N=$ between 10,000 to 1,000,000 with $P$ = 100 covariates, and additional simulations varying $P$ and looking at categorical data in the Appendix. Full details are in Appendix I.4.
Number of Batches
@@
With datasets of size $N=100,000$ and $N=200,000$, we compared the clustering performance when considering different numbers of batches. Further details are in Appendix I.5.
5.2 MNIST
We applied our model to the MNIST collection of $N=60000$ images of handwritten digits 0-9 (LeCun et al., 2011). As in Mansinghka et al. (2016), we downsampled each image to $16\times 16$ pixels using bilinear interpolation via TensorFlow, and represented this as a 256-dimensional binary vector. Furthermore, to reduce computation on irrelevant pixels (e.g. those in corners of the image) we removed all pixels with fewer than 100 non-zero values across the whole dataset, leaving 176 covariates . We split this into 6 batches of $N=10000$ and set $K=20$ as the maximum clusters in each batch. True labels for the MNIST digits were not used at any point in the unsupervised clustering.
@@ 5.3 Electronic Health Record (EHR) data
We analysed an anonymised EHR dataset derived from The Health Improvement Network (THIN) database that comprised records from 289,821 individuals in the United Kingdom over the age of 80. For 94 long-term health conditions, we recorded which conditions were present in each individual to create a $289,821\times 93$ binary matrix whose $(i,j)$-entry indicates whether individual $i$ has condition $j$ (encoded as a 1) or not (encoded as 0). Further details on data preprocessing are provided in Appendix L. The aim of the analysis is to identify clusters of individuals with common patterns of co-occurring conditions (multimorbidity).
6 Results
6.1 Simulation Study
6.1.1 Frequency of Merge/Delete
More detailed results for simulations, including an assessment of the quality of variable selection with noisy data, can also be found in Appendix J.1. The main conclusions of the simulation study are as follows:
Incorporating merge/delete enables faster convergence in terms of wall-clock time: Generally, MerDel models are faster to converge. Figure S1 in Appendix J.1 shows that early accepted merge/delete moves allow for substantial early increases in the ELBO. However, as described in Section 3.1, there is a trade-off that must be made - models with laps = 1, for example, get close to convergence very fast but then take more time to fully converge. The inclusion of frequent merge/delete moves at the start of the algorithm can also lead to the algorithm jumping to a worse local optimum, reflected in slightly lower ARIs in some cases.
Better ARIs are achieved with merge/delete: Figure 1 and Appendix J.1 show that (especially in Simulations 1.1, 1.2 and 1.3) using laps = {2, 5, 10} generally allows for models with higher ARIs than the variational model with no merge/delete moves.
Merge/delete more accurately estimates the number of clusters: The model with no merge/delete moves tends to vastly overestimate the number of true clusters in the data due to the existence of small unnecessary clusters. MerDel mitigates this, with Figure S2 in Appendix J.1 showing that the model often is able to find the exact true number. Some sporadic small clusters remain when implementing merge/delete moves every 10 laps in Simulation 1.3, indicating we could need more frequent moves as $N$ increases.
6.1.2 Global Merge Simulations
Results for selected simulations with $N$ ranging from 50,000 to 500,000 can be seen in Table 1, with further results in Appendix J.2. In our simulations, we compare to the “gold standard” setting in which the full dataset may be analysed without batching using MerDel with or without parallelisation (denoted by ‘par’ and ‘full’ in our results tables). We see from Table 1 that FedMerDel gave extremely good clustering results with median ARIs only marginally lower than those for ‘full’ and ‘par’.
The time taken was substantially faster for both parallelised MerDel and FedMerDel as expected via distributed computing, where MerDel on the full dataset became infeasible as our data sizes became larger due to the computational requirements of manipulating large matrices. However, as hypothesised in Section 3.2, we saw that FedMerDel was faster than paralellised MerDel, and became substantially more so as $N$ increased (Figure 2). Global merge times ranged from a mean of 0.29s(greedy)/0.39s(random) for $N=20,000$ to a mean of 14.8s(greedy)/18.2s(random) for $N=500,000$, showing that the time taken for the global merge was not computationally prohibitive.
Different batches of data for a dataset provided extremely consistent results, eg. the standard deviation of ARIs across 10 different shuffles of data in an $N=50,000$ simulation ranged from 0.0003 to 0.005 for the 20 simulated datasets. Notably, even if a cluster resides in only one batch of data, there is no requirement in our algorithm for this cluster to merge with clusters from other batches (if it does not improve the ELBO), allowing our algorithm to hold sound for data which is not identically distributed across partitions.
Figure 2: Plot comparing the mean time taken by MerDel with and without parallelisation (‘par’ and ‘full’), and FedMerDel as we vary $N$ with 5 or 10 batches/cores in ‘Global Merge Simulations’, with an approximate 95% confidence interval (mean $\pm$ 1.96 ×$\times\frac{s.d}{\sqrt{n}}$). Results for FedMerDel are with the random search.
Table 1: ‘Global Merge Simulations’ results. We report the median and lower/upper quantiles across all 10 ‘shuffles’/initialisations and all 10 synthetic datasets. FedMerDel results use random search; greedy search results are found in the Appendix, but differed little.
N Model Batches/Cores ARI Clusters Time (s)
50000 full 5 0.947 [0.941, 0.954] 12 [12, 12] 467 [414, 554]
par 5 0.948 [0.942, 0.954] 12 [12, 12] 176 [163, 196]
FedMerDel 5 0.945 [0.939, 0.948] 12 [12, 12] 114 [106, 123]
100000 full 5 0.955 [0.944, 0.961] 12 [12, 12] 956 [861, 1116]
par 5 0.954 [0.944, 0.961] 12 [12, 12] 338 [310, 380]
FedMerDel 5 0.954 [0.943, 0.960] 12 [12, 12] 260 [232, 287]
200000 par 10 0.952 [0.947, 0.958] 12 [12, 12] 482 [446, 536]
FedMerDel 10 0.951 [0.946, 0.957] 12.5 [12, 13] 270 [243, 305]
500000 par 10 0.950 [0.943, 0.951] 12 [12, 12] 1121 [1056, 1303]
FedMerDel 10 0.949 [0.942, 0.951] 12.5 [12, 13] 688 [646, 797]
6.1.3 Number of Batches
We saw that as we increased the number of batches used, although there was no statistically significant difference at the 5% level in the accuracy of the clustering as measured by ARI, we did see more clusters (Appendix J.3). Nevertheless, ARI values were all above 0.93 with a mean of 0.957 across all simulations.
6.2 MNIST
We ran FedMerDel on 5 different ‘shuffles’ of the data into batches with $K=20$ maximum clusters in each batch, and our resulting models after a global merge consisted of 27-33 clusters. It is unreasonable to expect to find 10 clusters exactly due to the variation of handwritten digits within numbers. The increased number of clusters was not an artefact of the global merging; a run of MerDel on the MNIST test set of $N=10,000$ digits with 35 maximum clusters found between 33-35 clusters where most merges/deletes were rejected. Other unsupervised clustering models in the literature on MNIST datasets also found a high number of clusters (Hughes & Sudderth, 2013; Ni et al., 2020).
Figure S12 depicts the distribution of true number labels within each cluster, and demonstrates subclustering within digits in the ‘best’ model out of our 5 clustering structures. This is chosen as the model with the highest ELBO, an approach used in other variational models (Ueda & Ghahramani, 2002; Hughes & Sudderth, 2013). We expected imperfect performance due to the reduction of the image to a 1D vector; every pixel is treated as independent of one another, and we lose all information from the 2D images pertaining to correlation between nearby pixels. We found that, for example, 4, 7, 9 were often seen together in clusters; these numbers have quite similar shapes and have a variety of shapes (Figure S11). On the other hand, FedMerDel performed particularly well in distinguishing 0, 1, 2 and 6 from other digits; 95.95% of 1’s and 92.46% of 6’s were classified into clusters predominantly made up of that digit. Further classification rates can be found in Appendix K.2, as well as a comparison to other scalable clustering methods.
Figure 3: Three examples of digits assigned to Clusters 5 and 14 depicting 6’s and 0’s respectively, and two generated digits from the variational distribution of each cluster.
Furthermore, we can use the posterior distribution of our model to probabilistically produce new digits from each cluster, showing the generative potential in variational Bayes. An example of generated 0’s and 6’s is seen in Figure 3, and further examples can be seen in Appendix K.3.
6.3 Electronic Health Record (EHR) data
@@
We divided the data into 14 batches of size 20,000 and 1 batch of size 9,821, and ran FedMerDel with $K=20$ maximum clusters within each batch. After global merging, we obtained 8 clusters, which are summarised in Figure 4. Each cluster can be characterised by particular conditions. From top to bottom in Figure 4, we have clusters of individuals with: (i) cancer; (ii) Atrial fibrilation (AF) & arrhythmia; (iii) peripheral vascular disease & aortic aneurism; (iv) stroke; (v) AF & arrhythmia and stroke; (vi) blindness; (vii) dementia; and (viii) “generally multimorbid” individuals.
@@
Figure 4: Results of EHR clustering. Columns are health conditions, and rows correspond to clusters. The shading indicates the proportion of individuals in each cluster that had each health condition, while row labels indicate the number of individuals in each cluster. To improve visualisation, only the most prevalent health conditions are shown (see Appendix L for further details and coding of health conditions).
7 Conclusion

### v2

@@ 5 Federated Variational Inference
In this section, we propose a scalable clustering algorithm for large categorical datasets, FedMerDel (Federated MerDel), which addresses the challenge of large $N$ and enables federated learning (FL).
@@
The second ‘global’ step then combines the local clustering structures. Local clusters are frozen, and we merge similar local clusters into ‘global’ clusters. This is conceptually similar to approaches used in SNOB (Zuanetti2018) and SIGN (Ni2019), but both use MCMC methods which are computationally costly; local clustering in SNOB takes over 4 hours for a batch of size $N=1000$ in the authors’ simulations. We utilise a novel version of the merge move in MerDel to implement a ‘global merge’. By freezing local clusters in the global merge, the proposed clustering structure and ‘global ELBO’ across all observations can be efficiently calculated by utilising stored parameter values. The variational framework allows us to accept or reject global merges in a principled manner. Importantly, the full dataset is not required in the global merge.
In contrast to the approach considered in Section 4.1, instead of optimising individual components of the existing classical variational algorithm, our approach is designed to present a scalable clustering model which avoids these bottlenecks entirely by adopting a distributed computing approach in which each data batch is analysed independently and in parallel before ‘global merge’ steps are used to obtain a (global) mixture model for the full dataset. This FL approach eliminates the need for frequent communication and synchronization during the main variational updates, while data remains decentralised.
@@ 5.1 Estimating Global Clusters
Assume our data has been split into $B$ batches and MerDel has been run on each batch, with $K_{b}$ local clusters in Batch $b$. Let $K=\sum_{b=1}^{B}K_{b}$ be the total number of local clusters. We begin with $\alpha^{\ast}$ and $\epsilon^{\ast}$ parameters fixed from the completed local clustering steps. In the global step, we first ‘combine’ all our existing variational parameters from each batch. For $q(\pi)$, we concatenate all our $K_{b}$-length $\alpha^{\ast}$ vectors to one $K$-length $\alpha^{\ast}$ vector. Each $q(\phi_{kj})$ is already independent for each $(k,j)$ pair. We create a new matrix of responsibilities $r_{nk}$ as an $N\times K$ block diagonal matrix with blocks of $N_{b}\times K_{b}$ responsibility matrices from each batch. $\sum_{k=1}^{K}r_{nk}=1$ still holds for all $n$.
Let clusters $k_{1}$ and $k_{2}$ be considered for a merge. As in the local merge, we reassign responsibilities to cluster $k_{1}$ via $r^{new}_{nk_{1}}=r_{nk_{1}}+r_{nk_{2}}$ for all n=1,,Nn=1,...,N. This can be viewed as summing columns $k_{1}$ and $k_{2}$ in our responsibility matrix. We then use the definitions of $\alpha^{\ast}$, $\epsilon^{\ast}$ (Equation (11)) as follows to update cluster-specific parameters:
$\alpha^{\ast new}_{k_{1}}$ $=\alpha_{0}+\sum_{n=1}^{N}(r_{nk_{1}}+r_{nk_{2}})=\alpha^{\ast}_{k_{1}}+\alpha^{\ast}_{k_{2}}-\alpha_{0}$ (12)
$\epsilon^{\ast new}_{k_{1}jl}$ $=\epsilon_{jl}+\sum_{n=1}^{N}\mathbb{I}(x_{nj}=l)(r_{nk_{1}}+r_{nk_{2}})=\epsilon^{\ast}_{k_{1}jl}+\epsilon^{\ast}_{k_{2}jl}-\epsilon_{j}$ (13)
This avoids calculating sums over $N$. To complete the merge, we remove all parameters associated with cluster $k_{2}$ - $\alpha^{\ast}_{k_{2}}$, $\epsilon^{\ast}_{k_{2}j}$ and $r_{nk_{2}}$ for all n=1,,Nn=1,...,N, j=1,,Pj=1,...,P. Unlike local merge moves, we do not perform additional E/M steps to allow observations to move in and out of the new cluster. This would require a whole-dataset update of responsibilities $r_{nk}$. The efficiency of the global merge hinges on local clusters being frozen. We propose global merge candidates using two approaches: greedy search and random search (see Appendix D.1 for details).
@@ 5.2 Efficient Global ELBO Calculation
The ELBO variational objective function allows us to accept or reject the global merge under a rigorous probabilistic framework. For the global merge to be accepted, the ELBO must be higher. As we now consider the full dataset, many terms in the overall ELBO function involve all observations and clusters. ‘Freezing’ local clusters allows us to efficiently calculate the majority of the ELBO terms and enable the model to handle large-scale data. Two sums involving $N$ in the full ELBO function can be written as a sum over smaller $K$, $P$, $L_{j}$ of sufficient statistics which can in turn be calculated as a linear function of previously calculated $\alpha^{\ast}$ and $\epsilon^{\ast}$: $T_{k}=\sum_{n=1}^{N}r_{nk}=\alpha^{\ast}_{k}-\alpha_{0}$ and $S_{kjl}=\sum_{n=1}^{N}r_{nk}\mathbb{I}(x_{nj}=l)=\epsilon^{\ast}_{kjl}-\epsilon_{jl}$. This vastly reduces the number of operations required.
The final term involving $N$ is an assignment entropy term, $\sum_{n=1}^{N}\sum_{k=1}^{K}r_{nk}\ln r_{nk}$, which is not linear in any previously calculated parameters. This is additive over $N$ and $K$, so we can calculate this value separately for each batch (in parallel if required) before summing results. However, in our greedy search, we never combine two clusters from the same batch, so $r_{nk_{1}}$ and/or $r_{nk_{2}}$ will always be zero for any merge, for all n=1,,Nn=1,...,N. Observation $n$ is only explained by clusters originating from some batches of data and a merge is only proposed from a new batch where $n$ is not explained by any clusters associated with that batch. This allows greedy search to be faster, although merges could be missed. See Appendix D.2 for full details of the global ELBO calculation.
Crucially, the global merge move allows us to seek a clustering structure for the full dataset without any scan over the full data matrix $X$. Data from separate entities remains decentralised, where only summary statistics from each batch are required to be shared, allowing for increased data security and privacy. These summary statistics represent cluster-specific weighted averages over batch-level populations, and do not directly expose individual-level data. We note that privacy risks could occur with clusters where these weights are almost all estimated to be close to 0 with a small number close to 1, but such cases can be easily detected and handled (eg. by withholding the cluster). Implementation details are in Appendix F. Parallelisation in local batches (Section 4.1) could be applied to FedMerDel for additional speed gains.
6 Experiments and Results
An additional study looking at unsupervised clustering of the MNIST dataset (MNIST) can be found in Appendix I.
@@ 6.1 Simulation Study
In this section, simulated data are binary, with the data generating mechanism detailed in the Appendix. Additional simulation studies comparing merge criteria, looking at categorical data with 3+ categories, and examining variable selection are in the Appendix.
6.1.1 Frequency of Merge/Delete Moves
@@
In Section 4, we described the trade-off between more or fewer merge/delete moves in MerDel. We ran simulation studies comparing different values of the parameter ‘laps’ representing the frequency of merge/delete moves in a run of MerDel, with laps $\in\{1,2,5,10\}$. We also compared to a model with just merge/delete moves after performing one ‘E’ and ‘M’ step after initialisation (‘laps = 0’), and the variational model with no merge/delete moves (‘laps = 10000’). We compared the wall-clock time taken for convergence, number of non-empty clusters found, and the Adjusted Rand Index (ARI) with the simulated ‘true’ labels. Appendix G.3 gives further details.
We saw that models incorporating merge/delete moves (laps $\in\{2,5,10\}$) consistently outperformed standard variational inference across multiple simulated datasets, achieving higher ARIs, more accurate cluster number estimation, and faster convergence (full results in Appendix H.1). Based on these results, we used laps =5=5 for subsequent experiments.
6.1.2 Global Merge Simulations
@@
We evaluated the clustering performance when splitting data into batches and performing a global merge as detailed in Section 5. In our simulations, we compared to the “gold standard” setting in which the full dataset may be analysed without splitting of the data using MerDel with or without parallelisation (denoted by ‘par’ and ‘full’ in our results tables). We looked at binary data with sizes varying between $N=$ between 10,000 to 1,000,000 with $P$ = 100 covariates.
For our simulated datasets, we generated data with known ground truth cluster structure, then initially randomly partitioned this into batches. We also looked at scenarios with heterogeneous data across partitions, where clusters are deliberately generated such that they may only appear in certain batches and not in others. We compared this to parallelised VI on the full data and compared to existing unsupervised federated and distributed algorithms such as k-FED (pmlr-v139-dennis21a), FedEM (NEURIPS2021_f740c8d9) and SNOB (Zuanetti2018). These represent the current state-of-the-art for unsupervised federated clustering, particularly for categorical data, in this nascent field. We also compared to some centralised learning methods as well as simulations varying $P$ and looking at categorical data in the Appendix. Full details are in Appendix G.4.
A full results table can be found in Appendix H.2, with results for heterogeneously distributed data shown in Table LABEL:globalmergeresults.
[…]
globalmergeresults Scenario Model Batches/Cores ARI Clusters Time (s) (a) par 10 0.945 [0.944, 0.951] 12 [12, 12] 119 [110, 131] FedMerDel 10 0.942 [0.936, 0.946] 13 [12, 13] 64.5 [61.5, 70.1] (b) par 5 0.955 [0.949, 0.956] 10 [10, 10] 169 [161, 192] FedMerDel 5 0.993 [0.992, 0.994] 10 [10, 10] 131 [127, 165] (c) par 5 0.950 [0.944, 0.950] 12 [12, 12.8] 69.5 [65.8, 77.3] FedMerDel 5 0.988 [0.985, 0.990] 13 [12, 13] 53.2 [48.3, 58.3]
Table 1: A subset of results for ‘Global Merge Simulations’, where data is distributed heterogeneously. For (a), (b), (c), the heterogeneous data scenarios are described in Appendix G.4. We report the median and lower/upper quantiles across all 10 ‘shuffles’/initialisations and all 10 synthetic datasets. FedMerDel results use random search.
When the data were divided into random batches, FedMerDel achieved clustering performance nearly equivalent to the “best possible” centralised baselines, with median ARIs only marginally below those of the full and parallel models (all above 0.9). While local batch-level ARIs were also high, the global merge step provides additional value by producing a unified model for the entire population. In heterogeneous scenarios, FedMerDel occasionally surpassed the full-data baseline (ARI $\approx$ 0.99). This occurs when batches contain fewer, well-separated clusters, making local clustering more accurate and improving the global merge outcome. Thus, FedMerDel accommodates statistical heterogeneity without forcing clusters to merge across batches, and can offer clustering-quality advantages beyond computational efficiency.
The time taken was substantially faster for both parallelised MerDel and FedMerDel as expected via distributed computing, where MerDel on the full dataset became infeasible as our data sizes became larger due to the computational requirements of manipulating large matrices. However, as hypothesised in Section 4.1, we saw that FedMerDel was faster than parallelised MerDel, and became substantially more so as $N$ increased (Figure LABEL:Nlines).
[…]
Nlines
Figure 1: Plot comparing the mean time taken by MerDel with and without parallelisation (‘par’ and ‘full’), and FedMerDel as we vary NN with 5 or 10 batches/cores in ‘Global Merge Simulations’, with an approximate 95% confidence interval (mean $\pm$ 1.96 $\times\frac{\text{s.d}}{\sqrt{n}}$). FedMerDel results use random search.
Global merge times ranged from a mean of 0.29s(greedy)/0.39s(random) for $N=20,000$ to a mean of 14.8s(greedy)/18.2s(random) for $N=500,000$, showing that the time taken for the global merge was not computationally prohibitive. Different batches of data for a dataset also provided consistent results, e.g. the standard deviation of ARIs across 10 different shuffles of data in an $N=50,000$ simulation ranged from 0.0003 to 0.005 for 20 datasets.
FedMerDel outperformed all comparison methods in terms of both accuracy and time efficiency. k-FED performed poorly (ARI between 0-0.2). FedEM involved frequent communication between clients and a central server, and took 1.5-4 hours per model. Local Gibbs sampling in parallel took 20-40 mins; FedMerDel took <<3 mins (Table LABEL:comparisontable). We also compared FedMerDel to non-federated clustering approaches including k-modes, hierarchical clustering, DBSCAN, and latent class analysis (LCA; Appendix H.2.1). While the heuristic algorithms were computationally efficient, they produced near-zero ARIs on our simulated datasets, indicating poor recovery of true clustering structures. LCA performed better (mean ARI $\approx$ 0.87 at $N=50,000$) but required multiple model fits and longer runtimes ($\sim$10 min per model).
[…]
comparisontable Sim FedEM Approx SNOB FedMerDel N=20000 10 cl: 0.847 [0.798, 0.865] 5 ba: 0.699 [0.695, 0.700] 5 ba: 0.920 [0.912, 0.933] 20 cl: 0.906 [0.849, 0.913] N=50000 25 cl: 0.839 [0.793, 0.879] 10 ba: 0.836 [0.814, 0.861] 10 ba: 0.951 [0.946, 0.954] Heterogeneous (c) 10 cl: 0.907 [0.902, 0.913] 5 ba: 0.502 [0.439, 0.632] 5 ba: 0.988 [0.985, 0.990]
Table 2: ARI results comparing unsupervised FL methods. k-FED is omitted due to poor performance. Where NN is stated, observations were assigned to partitions randomly; ‘heterogeneous (c)’ is described in Appendix G.4. ‘Cl’=clients, ‘ba’=batches.
6.1.3 Number of Batches
@@
With datasets of size $N=100,000$ and $N=200,000$, we compared the clustering performance when considering different numbers of batches (further details in Appendix G.5). As we increased the number of batches used, although there was no statistically significant difference at the 5% level in the clustering accuracy as measured by ARI, we did see more clusters (Appendix H.3). ARI values were all above 0.93.
@@ 6.2 Electronic Health Record (EHR) data
We analysed an anonymised EHR dataset derived from The Health Improvement Network (THIN) database that comprised primary-care records from 289,821 individuals in the United Kingdom over the age of 80. For 94 long-term health conditions, we recorded which conditions were present in each individual to create a $289,821\times 93$ binary matrix whose $(i,j)$-entry indicates whether individual $i$ has condition $j$ (encoded as a 1) or not (encoded as 0). Further details on data preprocessing are provided in Appendix J. The aim of the analysis is to identify clusters of individuals with common patterns of co-occurring conditions (multimorbidity) across a population, allowing us to discover clinically meaningful subgroups that support population health research.
@@
We divided the data into 14 batches of size 20,000 and 1 batch of size 9,821 using random partitioning, and ran FedMerDel with $K=20$ maximum clusters within each batch. After global merging, we obtained 12 clusters, which are summarised in Figure LABEL:THINsummary. Each cluster can be characterised by particular conditions. From top to bottom in Figure LABEL:THINsummary, we have clusters of individuals with: (i) cancer; (ii) Atrial fibrilation (AF) & arrhythmia; (iii) peripheral vascular disease & aortic aneurysm; (iv) stroke; (v) AF & arrhythmia and stroke; (vi) stroke and blindness; (vii) blindness; (viii) dementia; (ix) asthma and chronic obstructive pulmonary disease; (x)-(xii) “generally multimorbid” individuals.
To characterise these clusters and quantitatively validate their clinical relevance, we examined the mean number of conditions per person in clusters and cluster-level mortality statistics, which serve as an objective clinical outcome measure (see Appendix J.1). We found that our multimorbidity clusters were associated with different mortality levels, providing quantitative evidence that FedMerDel identified clinically relevant patient subgroups. For example, the stroke + AF/arrhythmia cluster has higher mortality than the stroke-only cluster, highlighting potential differences in clinical risk profiles. Of the “generally multimorbid” clusters, (x) sees a lower mean number of conditions per person and a lower mortality rate compared to the population average, (xi) is associated with psychiatric disorders and lower mortality, while cluster (xii) represents a high-mortality multimorbid group. While there may be further subclusters in cluster (x), we believe these do not emerge due to weak signals compared to core clusters. Future work using recursive clustering approaches could enable the discovery of such finer-grained structure.
[…]
THINsummary Refer to caption
@@
Figure 3: EHR clustering results with random search. Columns are health conditions, and rows correspond to clusters. Shading indicates proportion of individuals in each cluster with each condition; row labels show the number of individuals per cluster. To improve visualisation, only the most prevalent health conditions are shown. See Appendix J for further details, coding of health conditions and greedy search results.
To evaluate performance under realistic heterogeneity expected when combining data from different healthcare sites, we simulated heterogeneity in the THIN dataset (details and full results in Appendix J.2). When dementia patients were separated into distinct batches, the method revealed subclusters within the dementia group, supporting future work into recursive clustering. Using a Dirichlet distribution to simulate condition-based skews across batches, core clusters were retained, while smaller subclusters emerged, likely driven by batch-specific over-representation. Beyond computational efficiency, this federated approach offers valuable insights into cluster prevalence in batches - in real EHR settings, such patterns could identify region-specific multimorbidity clusters based on different sociodemographic characteristics of local populations. FedMerDel recovers meaningful population-level patterns even in the presence of inter-node variability.
7 Discussion